# A Survey of Vibe Coding with Large Language Models

### v1

@@ 5.3.2 Code-based Action Implementation
Contemporary systems implemented structured workflows through five-stage processes: planning, selection, parameter extraction, invocation, and result integration [100, 419, 74], following systematic patterns integrating generation with continuous mechanisms creating closed-loop systems where outcomes influence planning. Advanced workflows implement sophisticated error recovery creating robust pipelines, with environment interaction relying on multi-turn patterns following ReAct iteratively gathering information and taking actions [103, 74, 348].
@@
Code generation emerged as unified action paradigm consolidating diverse agent behaviors [243, 194, 420]. Executable code-based approaches [256] integrated Python interpreters enabling dynamic revision of prior actions upon observations through multi-turn interactions, achieving substantial improvements across API benchmarking compared to JSON-based alternatives. Repository-level systems [99, 243] extended capabilities from isolated snippets to complex projects with intricate dependencies, leveraging external tools for information retrieval, navigation, and implementation assistance. Multi-agent architectures [104] decomposed programming tasks through specialized coordination with programmer, test designer, and executor agents achieving superior performance while reducing computational overhead, while repository-scale evolution systems [421] demonstrated autonomous improvement through iterative planning across hundreds of files and thousands of lines, with evolved implementations outperforming human-designed solutions. Reinforcement learning approaches [165] enabled learning optimal strategies through policy optimization producing implementations mirroring human preferences and achieving state-of-the-art results while surpassing reference solutions, with agent frameworks [15] establishing comprehensive platforms enabling code writing, command-line operations, and multi-agent coordination supporting diverse implementations across multiple benchmarks. Beyond conventional synthesis tasks, code generation demonstrated versatility for structured knowledge applications, with KnowCoder pioneering Python class-based schema representations enabling unified information extraction through two-phase learning framework combining code pretraining and instruction tuning [422]. Building upon this foundation, KnowCoder-X subsequently addressed cross-lingual transfer limitations by standardizing multilingual schemas as Python classes, achieving substantial improvements over commercial systems through bilingual alignment instruction tuning [423]. Further advancing this paradigm, KnowCoder-V2 extended capabilities from extraction to deep analysis by bridging knowledge organization with analytical reasoning via unified code generation, thereby enabling systematic processing of complex knowledge analysis tasks [424].
@@ 5.4.1 Iterative Refinement
Effective code generation requires systematic refinement beyond single attempts [35, 116]. Large Language Models demonstrated significant potential in automated code generation [35]. However, accuracy remains limited particularly for complex tasks demanding comprehensive understanding [118]. This stems from inherent challenges in simultaneously comprehending natural language and producing correct code without automatic refinement, making single-attempt approaches insufficient [118, 108]. Researchers developed multi-round frameworks employing iterative refinement, significantly improving synthesis quality [105, 117, 425]. These enable LLMs to learn from errors through reflection mechanisms incorporating failed tests and using outcomes to enhance subsequent attempts [118]. The approaches evolved to incorporate feedback from multiple sources [109, 127]. Direct approaches integrate compiler feedback where code LLMs generate, execute, and improve code based on results [117]. Modern frameworks employ multi-agent architectures with coder and critic agents, where coder generates and critic analyzes providing feedback in iterative loops [272, 127, 104, 10]. Sophisticated systems implement specialized roles such as Reflection Agents, Thinking Agents, and Execution Agents [106].
@@ 5.4.2 Code Validation
Code validation evolved into sophisticated multi-layered approaches using dedicated critic agents conducting thorough analyses focusing on efficiency and functional correctness [426, 116, 127]. Automated testing provides quantitative metrics, categorized into using built-in benchmark test sets or LLM-generated test sets [426]. Advanced approaches leverage LLM-as-a-Judge frameworks, producing not only rating scores but also detailed feedback to repair incorrect code [427].
@@ 5.4.3 Intelligent Debugging
Self-reflection represents fundamental shift from one-shot generation to iterative refinement, involving prompting LLM to review earlier outputs identifying logical flaws [105, 116]. Research identified post-execution self-debugging, analyzing code after execution, struggling with bias [428]. In contrast, in-execution self-debugging examines intermediate states showing more promise [110, 429, 428]. Advanced frameworks employ search algorithms systematically exploring multiple debugging paths, enabling comprehensive repair capabilities [275]. Testing and debugging methodologies involve evaluating across multiple domains [430, 431, 432, 181, 425]. Industrial deployment shows promising results, with Meta’s TestGen-LLM achieving 75% building correctly, 57% passing reliably, and 25% increasing coverage [181]. The evolution toward multi-agent collaboration represents significant advancement mimicking real-world development practices [111]. Modern frameworks implement specialized architectures coordinating debugging with specialized agents as bug explainers, fault locators, and patch proposers [107]. Integration of multiple debugging and reflection techniques demonstrates exceptional improvements, with some reaching 98.2 on HumanEval employing debugging mechanisms decomposing failed code using Control Flow Graph analysis [111].
@@ 5.5.1 Collaboration Mechanisms
Complex software engineering tasks increasingly demand coordinated efforts beyond single-agent capabilities [124, 252]. Multi-agent systems represent significant shift from single-agent approaches to collaborative frameworks enabling multiple agents to collaborate by distributing responsibilities, facilitating scalable execution, improving resilience, and allowing dynamic adaptation [125, 124, 433, 127]. Communication serves as cornerstone of coordination, representing medium through which collective reasoning emerges [125], with LLMs enhancing communication, coordination, and decision-making enabling agents to interpret complex instructions [434, 435, 433, 10]. In software engineering contexts, LLM Multi-Agent Systems represent emerging approach where multiple agents are assigned specific roles to collaboratively execute complex development tasks, dividing responsibilities among specialized agents emulating real-world roles [436]. The field has developed several distinct communication architectures including four primary patterns: layered, decentralized, centralized, and shared message pools [119, 436]. Complementing these communication structures, role-based collaboration represents prominent strategy with LLM-based agents assuming specialized roles assigned sub-tasks to solve objectives [126, 252, 437], demonstrating considerable success with agents typically assigned roles based on expert knowledge implemented in frameworks like AutoGen and CrewAI using teams with different roles through task decomposition and specialization [120]. This approach categorizes into pipeline approaches where agents sequentially complete task parts, group discussion frameworks where agents communicate to reach agreements, or hybrid combinations [435, 438, 121, 437]. Specifically, shared message pool represents sophisticated communication approach serving as centralized hub where agents post, access, and interpret messages facilitating asynchronous reasoning. In MetaGPT, the pool operates as structured system enabling collaboration where agents publish and subscribe to typed messages, reducing communication overhead [436, 10].
@@ 5.5.2 Framework Implementations
Building upon these mechanisms, several frameworks demonstrate practical implementations of multi-agent collaboration. MetaGPT enables collaborations streamlining software engineering workflow through role-specific prompts establishing effective cooperation implementing standard operating procedures [368, 122, 252]. Similarly, ChatDev represents chat-powered framework where specialized agents are guided via chat chain modeling virtual software company communicating to autonomously generate software [10, 439, 122]. Additionally, CAMEL introduces framework enabling autonomous cooperation through role-playing, providing theoretical foundations for agent interaction. Multi-agent role-based collaboration has demonstrated consistent performance improvements leveraging collective intelligence [434]. MapCoder employs four specialized agents showcasing exceptional capabilities, achieving state-of-the-art results including 93.9% on HumanEval, 83.1% on MBPP, 22.0% on APPS, 28.5% on CodeContests, and 45.3% on xCodeEval [127]. MetaGPT demonstrated impressive accuracy through Standardized Operating Procedures [252, 121]. Beyond performance metrics, role-based systems offer significant cost advantages. ChatDev demonstrates how communicative agents facilitate seamless workflows, proving cost-effective and capable of proactively addressing errors, simplifying processes and improving code quality by mimicking human teamwork patterns [10]. These implementations collectively represent fundamental shift toward collaborative intelligence in software engineering [125].
@@ 6.1.1 Containerization Technologies
Containerization has become a fundamental component of the modern coding agent infrastructure. By encapsulating software, dependencies, and runtime configurations within portable images, containerization guarantees a consistent execution environment across diverse hardware and operating systems [440, 128, 129]. Containers rely on operating system–level virtualization to achieve lightweight isolation, rapid startup, and efficient resource utilization compared with full virtual machines [440, 130]. OS-level abstraction enables superior resource efficiency with quicker startup than virtual machines while providing agility, portability, reproducibility, modularity, and flexibility [129, 441, 128, 442, 443, 131, 444, 445, 446]. These properties are particularly valuable for coding agents, which often need to execute user-generated code repeatedly and securely while maintaining reproducibility and scalability [447, 448, 132, 133].
Docker remains the dominant technology for constructing and distributing container images. Its layered architecture, based on Dockerfiles and registries, allows for incremental builds and efficient updates [134, 449].In addition to Docker, lightweight virtualization frameworks such as Linux Containers (LXC/LXD) and specialized research sandboxes, including SandboxFusion and MPLSandbox, are increasingly employed in experimental systems that demand stronger isolation or multi-language support [450, 451, 135].
At the orchestration level, Kubernetes has become indispensable in managing containerized workloads. It assigns coding agent tasks into isolated pods, allocates computational quotas, and enforces automatic timeouts to prevent resource monopolization [136]. These mechanisms collectively ensure that agent-generated code can be executed in a controlled, reproducible, and efficient manner.
@@ 6.1.2 Security Isolation Mechanisms
Because large language models can generate arbitrary and potentially unsafe code, strict security isolation is essential. Sandbox-based systems provide the first line of defense by limiting the privileges of agent-executed code [11, 137, 138, 139, 452]. Tools such as gVisor intercept system calls to enforce fine-grained safety policies, while lightweight hypervisors and microkernels enhance separation between host and container [453]. Multi-layer sandboxing frameworks such as SAFE-LLM and NatiSand combine container isolation with dynamic monitoring to detect unsafe system behavior [137, 454, 455, 456].
Hardware-assisted isolation techniques further strengthen security. Intel PKRU memory protection and ARM TrustZone restrict unauthorized memory access and system control [457]. WebAssembly-based execution engines also provide deterministic and memory-safe environments for running untrusted agent code [458].
Contemporary coding agents typically embed execution backends that incorporate both static and dynamic policy enforcement. These backends restrict file I/O, network connections, and environment variables, while maintaining extensive logging to support traceability and auditability [459, 138]. Such hybrid security design enables continuous testing of agent-generated code without exposing the host system to unverified or malicious behaviors.
@@ 6.1.3 Cloud-based Execution Platforms
Cloud-based execution platforms extend the principles of containerization and sandboxing to distributed environments. They enable scalable execution across clusters, support dynamic resource allocation, and facilitate transparent workload scheduling [460, 461, 462]. Production deployments operate on 25,000 CPU core clusters organized into specialized pod configurations separating judging pods for evaluation from execution pods for code running with exclusive core assignment ensuring isolation and consistent performance, supporting AWS, Azure, and Google Cloud Platform with Kubernetes-based solutions enabling transparent remote execution [463].
Resource management involves containerized environments using Docker ensuring consistent execution across infrastructure configurations with rapid system reset capabilities [464, 465]. Security isolation relies on containerized sandbox environments creating multiple protection layers with Docker containers providing isolated file systems, process spaces, and network spaces implementing strict CPU and memory limits through Kubernetes management systems enforcing timeout strategies automatically terminating environments exceeding specified durations [136, 466, 467].
Runtime environments support diverse programming language ecosystems including C, C++, Java, Python, Rust, Go, C#, and PHP with execution pods handling specific compilation and runtime requirements enabling agents to work across different paradigms. Platforms use reproducible cloud compute instances providing consistent virtual hardware configurations deployed on standardized instances like Amazon EC2 with execution engines like JUDGE0 supporting over 60 languages through sandboxed Docker environments [466]. Modern implementations employ dedicated Google Cloud Platform servers with standardized configurations ensuring consistent execution eliminating interference between optimization runs [465]. Distributed frameworks implement two-tier pod design patterns separating execution from evaluation with large-scale implementations using 8,500 execution pods running individual code instances at concurrency 1 with exclusive core assignment while 2,000 judging pods handle complete evaluation workflows including correctness checking. Container orchestration follows microservices patterns where each agent and execution module operates in Kubernetes-managed containers with shared volumes and persistent volume claims facilitating data exchange while maintaining long-term storage of execution artifacts, logs, and results implementing secure REPL sandbox patterns governed by strict execution constraints [467].

### v2

@@ 5.3.2 Code-based Action Implementation
Contemporary systems implemented structured workflows through five-stage processes: planning, selection, parameter extraction, invocation, and result integration [srinivas2024retrieval, sapkota2025agents, yao2022react], following systematic patterns integrating generation with continuous mechanisms creating closed-loop systems where outcomes influence planning. Advanced workflows implement sophisticated error recovery creating robust pipelines, with environment interaction relying on multi-turn patterns following ReAct iteratively gathering information and taking actions [lin2025agent, yao2022react, huang2024levels].
@@
Code generation emerged as unified action paradigm consolidating diverse agent behaviors [dong2025survey, ishibashi2024self, zhao2025llm]. Executable code-based approaches [wang2024executable] integrated Python interpreters enabling dynamic revision of prior actions upon observations through multi-turn interactions, achieving substantial improvements across API benchmarking compared to JSON-based alternatives. Repository-level systems [zhang2024codeagent, dong2025survey] extended capabilities from isolated snippets to complex projects with intricate dependencies, leveraging external tools for information retrieval, navigation, and implementation assistance. Multi-agent architectures [huang2023agentcoder] decomposed programming tasks through specialized coordination with programmer, test designer, and executor agents achieving superior performance while reducing computational overhead, while repository-scale evolution systems [yu2025autonomous] demonstrated autonomous improvement through iterative planning across hundreds of files and thousands of lines, with evolved implementations outperforming human-designed solutions. Reinforcement learning approaches [gehring2024rlef] enabled learning optimal strategies through policy optimization producing implementations mirroring human preferences and achieving state-of-the-art results while surpassing reference solutions, with agent frameworks [wang2024openhands] establishing comprehensive platforms enabling code writing, command-line operations, and multi-agent coordination supporting diverse implementations across multiple benchmarks. Beyond conventional synthesis tasks, code generation demonstrated versatility for structured knowledge applications, with KnowCoder pioneering Python class-based schema representations enabling unified information extraction through two-phase learning framework combining code pretraining and instruction tuning [li2024knowcoder]. Building upon this foundation, KnowCoder-X subsequently addressed cross-lingual transfer limitations by standardizing multilingual schemas as Python classes, achieving substantial improvements over commercial systems through bilingual alignment instruction tuning [zuo2024knowcoder]. Further advancing this paradigm, KnowCoder-V2 extended capabilities from extraction to deep analysis by bridging knowledge organization with analytical reasoning via unified code generation, thereby enabling systematic processing of complex knowledge analysis tasks [li2025knowcoder].
@@ 5.4.1 Iterative Refinement
Effective code generation requires systematic refinement beyond single attempts [chen2021evaluating, jin2025reveal]. Large Language Models demonstrated significant potential in automated code generation [chen2021evaluating]. However, accuracy remains limited particularly for complex tasks demanding comprehensive understanding [jin2024rgd]. This stems from inherent challenges in simultaneously comprehending natural language and producing correct code without automatic refinement, making single-attempt approaches insufficient [jin2024rgd, jiang2023self]. Researchers developed multi-round frameworks employing iterative refinement, significantly improving synthesis quality [madaan2023self, bi2024iterative, grishina2025fully]. These enable LLMs to learn from errors through reflection mechanisms incorporating failed tests and using outcomes to enhance subsequent attempts [jin2024rgd]. The approaches evolved to incorporate feedback from multiple sources [zhang2024pair, islam2024mapcoder]. Direct approaches integrate compiler feedback where code LLMs generate, execute, and improve code based on results [bi2024iterative]. Modern frameworks employ multi-agent architectures with coder and critic agents, where coder generates and critic analyzes providing feedback in iterative loops [dong2023self, islam2024mapcoder, huang2023agentcoder, qian2023chatdev]. Sophisticated systems implement specialized roles such as Reflection Agents, Thinking Agents, and Execution Agents [kumar2025saarthi].
@@ 5.4.2 Code Validation
Code validation evolved into sophisticated multi-layered approaches using dedicated critic agents conducting thorough analyses focusing on efficiency and functional correctness [almorsi2024guided, jin2025reveal, islam2024mapcoder]. Automated testing provides quantitative metrics, categorized into using built-in benchmark test sets or LLM-generated test sets [almorsi2024guided]. Advanced approaches leverage LLM-as-a-Judge frameworks, producing not only rating scores but also detailed feedback to repair incorrect code [vo2025llm].
@@ 5.4.3 Intelligent Debugging
Self-reflection represents fundamental shift from one-shot generation to iterative refinement, involving prompting LLM to review earlier outputs identifying logical flaws [madaan2023self, jin2025reveal]. Research identified post-execution self-debugging, analyzing code after execution, struggling with bias [chen2025revisit]. In contrast, in-execution self-debugging examines intermediate states showing more promise [chen2023teaching, zhong2024ldb, chen2025revisit]. Advanced frameworks employ search algorithms systematically exploring multiple debugging paths, enabling comprehensive repair capabilities [goues2019automated]. Testing and debugging methodologies involve evaluating across multiple domains [haroon2025how, majdoub2024debugging, ribeiro2023large, alshahwan2024automated, grishina2025fully]. Industrial deployment shows promising results, with Meta’s TestGen-LLM achieving 75% building correctly, 57% passing reliably, and 25% increasing coverage [alshahwan2024automated]. The evolution toward multi-agent collaboration represents significant advancement mimicking real-world development practices [ashrafi2025enhancing]. Modern frameworks implement specialized architectures coordinating debugging with specialized agents as bug explainers, fault locators, and patch proposers [vinh2025repeton]. Integration of multiple debugging and reflection techniques demonstrates exceptional improvements, with some reaching 98.2 on HumanEval employing debugging mechanisms decomposing failed code using Control Flow Graph analysis [ashrafi2025enhancing].
@@ 5.5.1 Collaboration Mechanisms
Complex software engineering tasks increasingly demand coordinated efforts beyond single-agent capabilities [he2025llm, hong2024metagpt]. Multi-agent systems represent significant shift from single-agent approaches to collaborative frameworks enabling multiple agents to collaborate by distributing responsibilities, facilitating scalable execution, improving resilience, and allowing dynamic adaptation [tran2025multi, he2025llm, sarkar2025survey, islam2024mapcoder]. Communication serves as cornerstone of coordination, representing medium through which collective reasoning emerges [tran2025multi], with LLMs enhancing communication, coordination, and decision-making enabling agents to interpret complex instructions [du2023improving, liang2023encouraging, sarkar2025survey, qian2023chatdev]. In software engineering contexts, LLM Multi-Agent Systems represent emerging approach where multiple agents are assigned specific roles to collaboratively execute complex development tasks, dividing responsibilities among specialized agents emulating real-world roles [ha2025evaluating]. The field has developed several distinct communication architectures including four primary patterns: layered, decentralized, centralized, and shared message pools [yao2024comal, ha2025evaluating]. Complementing these communication structures, role-based collaboration represents prominent strategy with LLM-based agents assuming specialized roles assigned sub-tasks to solve objectives [li2024more, hong2024metagpt, bell2025future], demonstrating considerable success with agents typically assigned roles based on expert knowledge implemented in frameworks like AutoGen and CrewAI using teams with different roles through task decomposition and specialization [lou2025drf]. This approach categorizes into pipeline approaches where agents sequentially complete task parts, group discussion frameworks where agents communicate to reach agreements, or hybrid combinations [liang2023encouraging, gao2023large, mishra2025teammedagents, bell2025future]. Specifically, shared message pool represents sophisticated communication approach serving as centralized hub where agents post, access, and interpret messages facilitating asynchronous reasoning. In MetaGPT, the pool operates as structured system enabling collaboration where agents publish and subscribe to typed messages, reducing communication overhead [ha2025evaluating, qian2023chatdev].
@@ 5.5.2 Framework Implementations
Building upon these mechanisms, several frameworks demonstrate practical implementations of multi-agent collaboration. MetaGPT enables collaborations streamlining software engineering workflow through role-specific prompts establishing effective cooperation implementing standard operating procedures [lu2023towards, khanzadeh2025agentmesh, hong2024metagpt]. Similarly, ChatDev represents chat-powered framework where specialized agents are guided via chat chain modeling virtual software company communicating to autonomously generate software [qian2023chatdev, mou2024from, khanzadeh2025agentmesh]. Additionally, CAMEL introduces framework enabling autonomous cooperation through role-playing, providing theoretical foundations for agent interaction. Multi-agent role-based collaboration has demonstrated consistent performance improvements leveraging collective intelligence [du2023improving]. MapCoder employs four specialized agents showcasing exceptional capabilities, achieving state-of-the-art results including 93.9% on HumanEval, 83.1% on MBPP, 22.0% on APPS, 28.5% on CodeContests, and 45.3% on xCodeEval [islam2024mapcoder]. MetaGPT demonstrated impressive accuracy through Standardized Operating Procedures [hong2024metagpt, mishra2025teammedagents]. Beyond performance metrics, role-based systems offer significant cost advantages. ChatDev demonstrates how communicative agents facilitate seamless workflows, proving cost-effective and capable of proactively addressing errors, simplifying processes and improving code quality by mimicking human teamwork patterns [qian2023chatdev]. These implementations collectively represent fundamental shift toward collaborative intelligence in software engineering [tran2025multi].
@@ 6.1.1 Containerization Technologies
Containerization has become a fundamental component of the modern coding agent infrastructure. By encapsulating software, dependencies, and runtime configurations within portable images, containerization guarantees a consistent execution environment across diverse hardware and operating systems [ferro2022containerization, santoro2023virtual, lee2023kunerva]. Containers rely on operating system–level virtualization to achieve lightweight isolation, rapid startup, and efficient resource utilization compared with full virtual machines [ferro2022containerization, lounissi2025funda]. OS-level abstraction enables superior resource efficiency with quicker startup than virtual machines while providing agility, portability, reproducibility, modularity, and flexibility [lee2023kunerva, sultan2019container, santoro2023virtual, pahl2019cloud, romanov2024optimizing, baresi2020cocos, amaral2015performance, morabito2017virtualization, pahl2015containers]. These properties are particularly valuable for coding agents, which often need to execute user-generated code repeatedly and securely while maintaining reproducibility and scalability [moulton2024confronting, boettiger2014introduction, adamidi2025virtual, kurtzer2017singularity].
Docker remains the dominant technology for constructing and distributing container images. Its layered architecture, based on Dockerfiles and registries, allows for incremental builds and efficient updates [zhou2022drive, sarishma2021systematic].In addition to Docker, lightweight virtualization frameworks such as Linux Containers (LXC/LXD) and specialized research sandboxes, including SandboxFusion and MPLSandbox, are increasingly employed in experimental systems that demand stronger isolation or multi-language support [baresi2023qualitative, miller2020towards, cassano2023multipl].
At the orchestration level, Kubernetes has become indispensable in managing containerized workloads. It assigns coding agent tasks into isolated pods, allocates computational quotas, and enforces automatic timeouts to prevent resource monopolization [su2024large]. These mechanisms collectively ensure that agent-generated code can be executed in a controlled, reproducible, and efficient manner.
@@ 6.1.2 Security Isolation Mechanisms
Because large language models can generate arbitrary and potentially unsafe code, strict security isolation is essential. Sandbox-based systems provide the first line of defense by limiting the privileges of agent-executed code [jimenez2023swe, rabin2025sandboxeval, nunez2024autosafecoder, gajbhiye2024secure, bahadur2025securing]. Tools such as gVisor intercept system calls to enforce fine-grained safety policies, while lightweight hypervisors and microkernels enhance separation between host and container [rokon2020sourcefinder]. Multi-layer sandboxing frameworks such as SAFE-LLM and NatiSand combine container isolation with dynamic monitoring to detect unsafe system behavior [rabin2025sandboxeval, liu2023your, siddiq2023sallm, abbadini2023natisand].
Hardware-assisted isolation techniques further strengthen security. Intel PKRU memory protection and ARM TrustZone restrict unauthorized memory access and system control [kirth2022pkru]. WebAssembly-based execution engines also provide deterministic and memory-safe environments for running untrusted agent code [kanellopoulos2024virtuoso].
Contemporary coding agents typically embed execution backends that incorporate both static and dynamic policy enforcement. These backends restrict file I/O, network connections, and environment variables, while maintaining extensive logging to support traceability and auditability [wang2025agentspeccustomizableruntimeenforcement, nunez2024autosafecoder]. Such hybrid security design enables continuous testing of agent-generated code without exposing the host system to unverified or malicious behaviors.
@@ 6.1.3 Cloud-based Execution Platforms
Cloud-based execution platforms extend the principles of containerization and sandboxing to distributed environments. They enable scalable execution across clusters, support dynamic resource allocation, and facilitate transparent workload scheduling [feilhauer2016def, abolfazli2013cloud, bhm2022cloud]. Production deployments operate on 25,000 CPU core clusters organized into specialized pod configurations separating judging pods for evaluation from execution pods for code running with exclusive core assignment ensuring isolation and consistent performance, supporting AWS, Azure, and Google Cloud Platform with Kubernetes-based solutions enabling transparent remote execution [tagliabue2023reasonable].
Resource management involves containerized environments using Docker ensuring consistent execution across infrastructure configurations with rapid system reset capabilities [vijayvargiya2025openagentsafety, gong2025tuning]. Security isolation relies on containerized sandbox environments creating multiple protection layers with Docker containers providing isolated file systems, process spaces, and network spaces implementing strict CPU and memory limits through Kubernetes management systems enforcing timeout strategies automatically terminating environments exceeding specified durations [su2024large, waghjale2024ecco, ardebili2025kubeintellect].
Runtime environments support diverse programming language ecosystems including C, C++, Java, Python, Rust, Go, C#, and PHP with execution pods handling specific compilation and runtime requirements enabling agents to work across different paradigms. Platforms use reproducible cloud compute instances providing consistent virtual hardware configurations deployed on standardized instances like Amazon EC2 with execution engines like JUDGE0 supporting over 60 languages through sandboxed Docker environments [waghjale2024ecco]. Modern implementations employ dedicated Google Cloud Platform servers with standardized configurations ensuring consistent execution eliminating interference between optimization runs [gong2025tuning]. Distributed frameworks implement two-tier pod design patterns separating execution from evaluation with large-scale implementations using 8,500 execution pods running individual code instances at concurrency 1 with exclusive core assignment while 2,000 judging pods handle complete evaluation workflows including correctness checking. Container orchestration follows microservices patterns where each agent and execution module operates in Kubernetes-managed containers with shared volumes and persistent volume claims facilitating data exchange while maintaining long-term storage of execution artifacts, logs, and results implementing secure REPL sandbox patterns governed by strict execution constraints [ardebili2025kubeintellect].